OpenStudy (anonymous):

Help me to condense an equation! Thanks! (:

OpenStudy (anonymous):

\[3(\log_{3}14-\log_{3}7)+ 2 \log_{3}5\]

OpenStudy (anonymous):

\[\log(14)-\log(7)=\log(\frac{14}{7})=\log(2)\] is a start

OpenStudy (anonymous):

then \[3\log(2)=\log(2^3)=\log(8)\]

OpenStudy (anonymous):

and finally \[\log(8)=2\log(5)\log(8)+\log(5^2)=\log(8\times 25)=\log(200)\]

jimthompson5910 (jim_thompson5910):

Made a typo, but I fixed it \[\Large 3\left(\log_{3}(14)-\log_{3}(7)\right)+ 2 \log_{3}(5)\] \[\Large 3\log_{3}\left(\frac{14}{7}\right)+ 2 \log_{3}(5)\] \[\Large 3\log_{3}\left(2\right)+ 2 \log_{3}(5)\] \[\Large \log_{3}\left(2^3\right)+ \log_{3}(5^2)\] \[\Large \log_{3}\left(8\right)+ \log_{3}(25)\] \[\Large \log_{3}(8*25)\] \[\Large \log_{3}(200)\] So \[\Large 3\left(\log_{3}(14)-\log_{3}(7)\right)+ 2 \log_{3}(5)=\log_{3}(200)\]